OpenStudy (anonymous):

Hi can anyone help me with a statistics problem: Here it is: You have a student, Martina, entering your school from another state. You need to know which math class you should place her in. The scores you normally use are SAT scores (mean=500 and standard deviation=100). The rules you use for placement are: 700-800 Calculus 500-699 Trigonometry 300-499 Geometry 100-299 Algebra However, this student did not take the SAT, but instead took the AAT. Scores for the AAT are reported with a mean of 80 and a standard deviation of 8. Martina received a raw score of 84 on the math portio

OpenStudy (campbell_st):

find the z score - or the standard score) for the AAT test \[z = (x - \mu)/\sigma\] \[z = (84 - 80)/8\] z = 0.5 so Matha scored 1/2 a standard deviation above the mean the equivalent for the SAT would be \[x = \mu + z \times \sigma\] her score would be 500 + 0.5 x 100 Martina would have an equivalent score of 550

OpenStudy (anonymous):

Thank you so much!